OpenStudy (anonymous):

The midpoint of UV is (5,-11). The coordinate of one endpoint are U(3,5). Find the coordinates of endpoint V .

OpenStudy (anonymous):

I'll fan and medal !!!

OpenStudy (disqin):

Use relation \[\frac{x_1+x_2}{2}=x\] and \[\frac{y_1+y_2}{2}=y\] Use \((x,y)=(5, -11)\) and \((x_1, y_1)=(3, 5)\) and solve for \(x_2\) and \(y_2\) like a boss B-)

OpenStudy (anonymous):

thank you so much, i'll let you know how it goes !